Circular No. 8758

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2006ha–2006iu Forty-seven additional apparent supernovae have been announced on

CBETs (those reported as brighter than mag 18.5 are tabulated below), including 25 supernovae found on g, r, and i images and reported by J. Frieman on behalf of the SDSS II collaboration (most in the magnitude range g = 20–23; 2006hc–2006hm, 2006ho–2006ib; details on CBETs 650, 654, 656, and 657), sixteen supernovae found by the “Nearby Supernova Factory” (NSF) collaboration (2006ic–2006ir; details on CBET 658); and the following six apparent supernovae discovered on unfiltered CCD images: 2006ha, 2006hb, 2006is, 2006it, and 2006iu by the Lick Observatory Supernova Search (via N. Joubert, R. Mostardi, R. R. Prasad, N. Lee, M. Baek, and W. Li; cf. IAUC 8754); and 2006hn by A. Sehgal, R. Gagliano, and T. Puckett (cf. IAUC 8754; 0.50-m reflector at Osoyoos, BC).

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006ha Sept. 27.24 22^{h}58^{m}34.^{s}30 +15^{o}10′25′′.8 18.4 0′′.6 W, 3′′.2 N
2006hb Sept. 27.52 5 02 01.28 −21 07 55.1 15.7 8′′.6 E, 18′′.2 N
2006hn Sept. 28.46 11 07 18.67 +76 41 49.8 15.6 7′′.6 N
2006im Sept. 14.3 22 07 58.11 −23 05 49.9 18.3  —
2006ir Sept. 23.3 23 04 35.68 + 7 36 21.5 16.9  —
2006is Sept. 18.54 5 17 34.37 −23 46 54.2 17.0 2′′.1 W, 4′′.9 N
2006it Oct. 1.17 20 43 51.38 +12 30 30.9 17.6 34′′.4 W, 12′′.0 S
2006iu Oct. 1.51 6 36 41.13 +37 36 05.0 17.9 0′′.6 E, 2′′.0 N

Additional magnitudes for 2006ha in IC 1461: Sept. 10.24 UT, [19.0; 19.25, [19.2; 26.25, 18.9. Additional magnitudes for 2006hb in MCG –04-12-34: Feb. 24.12, [19.0; Sept. 11.51, 18.0. SN 2006hb is a subluminous type-Ia supernova, near maximum brightness at discovery (details on CBET 652). Additional magnitudes for 2006hn in UGC 6154: June 8, [19.5 (Puckett); Sept. 29.21, 15.5 (Sehgal). Of the SDSS and NSF discoveries, eight are type-II supernovae (2006ho, 2006hz, 2006ic, 2006ih, 2006ii, 2006ij, 2006im, 2006in), two are type-Ib/c supernovae (2006ip, 2006ir), and the rest appear to be type-Ia supernovae. Additional magnitudes for 2006is: Feb. 4.11, [19.0; Sept. 19.54, 16.8; 28.51, 16.4. There are several faint galaxies visible on the Digitized Sky Survey images in the neighborhood of 2006is, which is a type-Ia supernova around maximum light on Sept. 30 (details on CBET 659). Additional magnitudes for 2006it in NGC 6956: Sept. 26.19, [19.2; Oct. 7.24, 17.5. SN 2006it is a type-IIP supernova (details on CBET 665). Additional magnitudes for 2006iu: Feb. 24.11, [19.0; Oct. 8.53, 18.1.

2006 October 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green